OpenStudy (anonymous):

Can anyone help me solve this problem -3-2x=-13

OpenStudy (anonymous):

-3-2X=-13 +3 +3 ___________________ -2X=-10 DIVIDE -2X TO BOTH SIDES= X=5

OpenStudy (anonymous):

DO U UNDERSTAND?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

AWESOME:)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

NO PROBLEM.:)